OpenStudy (shelby1290):

The function P(x)=-30x^2+360+785 models the profit, P(x), earned by a theatre owner on the basis of a ticket price, x. Both the profit and ticket price are in dollars. What is the maximum profit, and how much should the tickets cost?

imqwerty (imqwerty):

ok :) do u know differentiation?

OpenStudy (shelby1290):

never heard of it

OpenStudy (shelby1290):

this is not calculus math

imqwerty (imqwerty):

oh okk :) we can do it the other way in this way you have to use smart vision e.e ready?

OpenStudy (shelby1290):

@imqwerty sure, what's the second way of doing this?

imqwerty (imqwerty):

ok note this-\[p(x)=-30x^2 +360+785\] i'll tell step by step x^2 will always be positive do u agree?

OpenStudy (shelby1290):

yes

imqwerty (imqwerty):

ok and -30x^2 will also be negative do you agree?

OpenStudy (shelby1290):

yes

imqwerty (imqwerty):

ok more the x more will be the value of -30x^2 agreed?

OpenStudy (shelby1290):

What do you mean by "more the x more the value of -30x^2" ?

OpenStudy (shelby1290):

@imqwerty ^

OpenStudy (shelby1290):

I don't understand your question

imqwerty (imqwerty):

ok try putting x=1 and x=2 then find the value of -30x^2 that u get

OpenStudy (shelby1290):

am i putting x=1 into this equation? : p(x)=−30x2+360+785

imqwerty (imqwerty):

no jst put x=1 and then x=2 in -30x^2 and tell what u get

OpenStudy (shelby1290):

-30x^2 -30(1)^2 =-30 -30(2)^2 =-120

imqwerty (imqwerty):

yes can you conclude that more the value of x lesser wil -30x^2 become note that -120 is lesser than -30

OpenStudy (shelby1290):

oh okay i see

OpenStudy (shelby1290):

How do i find the maximum profit?

OpenStudy (welshfella):

are you sure the middle term is 360 - not 360x ?

OpenStudy (shelby1290):

@welshfella Yes it is 360x , not negative 360x

imqwerty (imqwerty):

ok now see this -\[p(x)=-30x^2+360+785\] here only negative term is -30x^2 the more its negative the more it will decrease the profit its max value can be 0 when x=0 so x=0 damn it its 360x!

OpenStudy (welshfella):

with 360 it would make the function a parabola opeeing downwards with a maximum turning point

OpenStudy (shelby1290):

@imqwerty okay, so what must I solve for? I'm a bit confused right now @welshfella mhm...

OpenStudy (welshfella):

wth 360x i meant to say

OpenStudy (welshfella):

well to find the maximum by completing the square and converting the equation to vertex form

OpenStudy (welshfella):

or you could use -b/2a which gives the x coordinate value when the function is a maximum

OpenStudy (mathmale):

@shelby1290: P(x)=-30x^2+360+785 is a quadratic function. Because the sign of the greatest power of x is negative, the graph of this function opens DOWN. We need only find the x value at which the profit is a maximum. Yes, finding x = -b/(2a) will give you that x-value. Graphing P(x)=-30x^2+360+785 may help you to visualize what is happening. Please identify coefficients a and b in P(x)=-30x^2+360+785.

OpenStudy (shelby1290):

@welshfella alright so like this? P(x)=(-30x^2+360x)+785 =-30(x^2+360x)+785 =-30(x^2-12x)+785 =-30(x^2-12x+36-36)+785 =-30(x^2-12x+36)-23550+785 =-30(x^2-12x+36)-22765 =-30(x-12)(x-12)-22765 =-30(x-12)^2-22765

OpenStudy (shelby1290):

@mathmale -b/2a = -(360)/2(-30) = -360/-60 =6

OpenStudy (mathmale):

Then calculate x=-b/(2a). Once you have that, calculate P(x). You've done a lot of work to complete the square. Can you explain the significance of your results? If your calculations are correct, your work gives you both the ticket cost for max profit and the actual max. profit.

OpenStudy (mathmale):

Here you've found that x=6 (6 tickets). But when you completed the square, you found that x=12. You're on the right track, but both of your results (x=6, x=12) can't be right. If you have a graphing calculator, graph P(x)=-30x^2+360+785 and determine from the graph the x value at which the graph has a max.

OpenStudy (shelby1290):

I think I made a mistake because the maximum profit is $1865

OpenStudy (shelby1290):

I don't have a graphing calculator with me

OpenStudy (mathmale):

Again, one result or the other (x=6, x=12) MAY be correct, but not both. Besides graphing P(x)=-30x^2+360+785, what other ways have you for finding the max. of P(x)?

OpenStudy (mathmale):

Looking at P(x)=-30x^2+360x+785, I see that a = -30 and b=360. These values disagree with yours. Can you explain why?

OpenStudy (shelby1290):

I've only learned how to find the max by finding the vertex, which is by completing the square

OpenStudy (mathmale):

OK, then. Let's look at the given function again: P(x)=-30x^2+360x+785. Before we move on, please double check to ensure that this equation is the correct one.

OpenStudy (shelby1290):

OH i see where i made a mistake now

OpenStudy (mathmale):

What mistake was that?

OpenStudy (shelby1290):

=-30(x^2+360x)+785 =-30(x^2-12x)+785 =-30(x^2-12x+36-36)+785 =-30(x^2-12x+36)+1080+785 =-30(x^2-12x+36)+1865 =-30(x-12)(x-12)+1865 =-30(x-12)^2+1865 I didn't multiply -36 by -30 to get 1080 and then 1080 +785 =1865

OpenStudy (shelby1290):

$1865 is the max profit

OpenStudy (mathmale):

Think you can fix this yourself? How could I help you further? Are you done now, or have y ou more work to do on this particular problem?

OpenStudy (welshfella):

yes - good work

OpenStudy (shelby1290):

I believe that's all I have to do because the question is only asking for the max profit and how much the tickets should cost

OpenStudy (mathmale):

Good for y ou! See you again. Bye.

hero (hero):

@Shelby1290 have you checked your messages?